1.  Consider an industry in the U.S. facing aggregate (inverse) demand function:

p(y) = 1050 – 5y


The industry is currently in long run equilibrium. The market price is $225 and there are n = 11 firms producing. Each firm’s variable cost is:


cv(y) = [1]  y3


a.  What is each firm’s fixed cost?



1
Expert's answer
2020-11-25T12:21:06-0500

let the fixed cost be "F"

total cost = "TC(y) = F + \frac{1}{3}\times y^{3}"

marginal cost "= MC(y) = \frac{d TC(y)}{dy} = y^{2}"

average cost "= \frac{TC(y)}{y} = \frac{F}{y}+\frac{1}{3}\times y^{2}"

set "MC(y) =AC(y)" to determine the output level at which "AC(y)" is minimized

"y^{2} = \frac{F}{y}+\frac{1}{3}\times y^{2}"

"\frac{2}{3}y^{2} = \frac{F}{y}"

"y^{3} = \frac{3}{2}F"

"y= (1.5F)^{\frac{1}{3}}"

so "MC" at this output is given by "MC(y*)" "=(1.5F)^{\frac{2}{3}}"

set "MC(y*) =" long run price

"(1.5F)^{\frac{2}{3}}= 225"

"1.5F=(225)^{\frac{3}{2}}=3375"

"F=\$2250" (fixed cost)